# Alternative Approaches for Computing Highest-Density Regions

### v1

@@ 3.1 $\boldsymbol{k}$NN-Euclidean distance
This corresponds to the sum of all Euclidean distances of a given point $x$ from its $k$-nearest neighbors. The concept of closeness (“nearest”) is defined according to the Euclidean metric or $L^{2}$-norm denoted by $\lVert\cdot\rVert_{2}$.
Definition 3 (kNN-Euclidean distance)
@@
where $x_{(i)}$ denotes the $i$-th observation in the reordered data such that $\lVert x-x_{(1)}\rVert_{2}\leq\dots\leq\lVert x-x_{(i)}\rVert_{2}\leq\dots\lVert x-x_{(n)}\rVert_{2}$.
$M_{1}(x,s_{n},k)$ is entirely based on a metric distance, that is, the Euclidean metric, and represents a sparsity measure. The property follows from the convergence in probability of the $k$-nearest neighbor density estimator (Silverman,, 1996). Similar distances have been used in the literature for nonparametric classification and regression, within the so-called $k$-nearest neighbors ($k$NN) algorithm. Notably, the distance to the nearest neighbors can also be seen as a local density estimate (Loftsgaarden and Quesenberry,, 1965; Silverman,, 1996) and as a special case of a variable-bandwidth KDE with a uniform kernel (Terrell and Scott,, 1992). A large $k$NN distance indicates that the density is usually small and vice versa. Furthermore, by ranking each data point according to the distance from its $k$-nearest neighbors, this measure can also be used as an outlier score in anomaly detection (Ramaswamy et al.,, 2000).
Compared to other methods, the $k$NN approach has several advantages such as: (i) being purely nonparametric, hence able to flexibly adapt to any continuous distribution; (ii) having a reasonable time complexity; (iii) depending on a unique hyperparameter $k$, whose tuning is relatively simple. The choice of $k$ should be made based on the sample data. Generally, higher values of $k$ reduce the effect of noise; however, they underfit the model–making boundaries between classes less distinct–and are computationally more expensive, especially for large $d$. A general rule of thumb in classification is $k=\lfloor\sqrt{n}\rfloor$, with $n$ the number of samples in the dataset. Further investigation of the role of $k$ in different settings and how it relates to the missclassification error is given in the Supplementary Material B and in (Meeker et al.,, 2017).
@@ 3.2 $\boldsymbol{k}$NN-CDF distance
Although the Euclidean distance and other general notions of measures (including concepts like area or volume) can certainly be useful for capturing relevant insights on the topology of a given set, they may be equipped with a probability measure to better resemble the notion of density. Let $\mathbb{P}$ be a probability measure and denote by $F$ its associated CDF and by $F_{i}$ the CDF of the $i$-th marginal.
@@
The distance $d_{\mathbb{P}}$ can be interpreted as the composition of the (nonlinear) transformation $F\!:\mathbb{R}^{d}\to[0,1]$ and the computation of the ordinary Euclidean distance. It fulfills all the properties of being a proper metric. Furthermore, it has the nice property that the distance between two points is proportional to the probability contained between the two points. However, as defined in Eq. (3), it presents some key limitations for inferring the underlying density and estimating HDRs. Consider for simplicity the univariate Gaussian mixture model illustrated in Figure 3 and the set of points $x_{1}=-2$, $x_{2}=-1.6$, $x_{3}=-0.72$, $x_{4}=0.5$. When evaluating the CDF distance between $x_{2}$ and all other points, we have that $d_{\mathbb{P}}(x_{2},x_{1})\approx d_{\mathbb{P}}(x_{2},x_{3})$, but there is little to say about the density around $x_{1}$ and $x_{3}$, apart from realizing that they are clearly different. We could gain more information if we were to compute the CDF distance from the global mode $x_{1}$, assuming we have this information, noticing that the higher $d_{\mathbb{P}}(x_{1},x_{i})$, for all $i$, the lower the density of $x_{i}$. However, this is no longer verified for multimodal densities. In fact, we have $d_{\mathbb{P}}(x_{1},x_{4})>d_{\mathbb{P}}(x_{1},x_{3})$, but $f(x_{4})>f(x_{3})$, contrasting with the definition of a neighborhood measure.
@@
Motivated by these limitations and inspired by the idea of the $k$NN distance in Definition 3, we propose a new variant consisting of the sum of CDF-distances between a point $x$ and its $k$-nearest neighbors. Here, “nearest” should again be understood according to the Euclidean distance. The rationale is simple: the higher the CDF distances between a point and its neighbors, the higher we expect the density to be for that point. Since the $k$ neighbors of each data point will be at different (Euclidean) distances compared to the $k$ neighbors of the other data points, we want to properly scale or weight the CDF distances according to this information. The proposed measure is reported in Definition 4.
Definition 4 ($\boldsymbol{k}$NN-CDF distance)
@@
Property 1 (Semimetric)
Consider the $M_{2}$ measure as defined in Definition 4. It can be easily verified that, for any $k\in[1,n]$, $M_{2}(x,x_{(k)},k)\!:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a semimetric. In fact, for any set $x,x_{(k)}\in\mathbb{R}^{d}$, we have that:
@@
The triangular inequality, stating that $M_{2}(x,x_{(k)},1)\leq M_{2}(x,x_{(k^{\prime})},1)+M_{2}(x_{(k^{\prime})},x_{(k)},1)$, for all $x,x_{(k)},x_{(k^{\prime})}\in\mathbb{R}^{d}$, does not hold. It is enough to consider a simple counterexample such as: $d=1$, $F$ the CDF of a standard normal distribution, and three points $x=1<x_{(k)}=2<x_{(k^{\prime})}=3$.
@@ 3.3 $\boldsymbol{\epsilon}$-neighborhood multivariate CDF distance
The measure $M_{2}(x,s_{n},k)$ introduced in Section 3.2 has several advantages; for example: (i) it takes into account the probabilistic information in the data, (ii) it is based on marginal CDFs which may be easy to estimate, and (iii) it has convenient computational times. However, using only the marginal CDFs may compromise their validity whenever the marginal components of a multivariate random variable $X\in\mathbb{R}^{d}$ show a dependence structure. In that case, the use of a multivariate CDF may be more appropriate. If the multivariate CDF is unknown, we could use its empirical counterpart. Notice, however, that this estimation part may compromise computational efficiency, especially in high dimensions and for large values of $k$. We therefore focus on an $\epsilon$-neighborhood version, which reduces the number of operations from $k$ to $1$, once $\epsilon$ is determined.
Definition 5 ($\epsilon$-neighborhood multivariate CDF distance)
@@
$\quad-\mathbb{P}(X_{1}\leq x_{1}+\epsilon_{1},X_{2}\leq x_{2}-\epsilon_{2})-\mathbb{P}(X_{1}<x_{1}-\epsilon_{1},X_{2}<x_{2}-\epsilon_{2})$
@@
Property 2 (Density equivalence)
@@ 3.4 Copula-based measures
In summary, we can decompose every $d$-dimensional CDF $F$ or PDF $f$ into a composition of their marginal distribution functions and a $d$-copula. This allows us to redefine both $M_{0}$ in Eq. (2) and $M_{3}$ in Definition 5 in terms of their copula representation. For example, in the case of the $M_{3}$ measure, we may construct its copula-based alternative given in Definition 6.
Definition 6 ($\epsilon$-neighborhood copula-based CDF distance)
@@
[…]
@@
The main advantage of this representation over the one involving the joint CDF is that the estimation of the multivariate distribution–when unknown–is performed through the estimation of the (univariate) marginals, evading thus the curse of dimensionality (see e.g., Nagler and Czado,, 2016). Furthermore, copulae offer a flexible framework that captures complex dependency structures while offering direct control over the marginals.
@@ 4.1 Simulation setup
We now evaluate and compare the proposed measures in an extensive number of simulated bivariate settings that vary according to the complexity of the data, as defined below.
@@
    We also vary the sample size, with $n\in\{50,100,500,1000\}$.
@@ 4.2 Evaluated measures
This represents the $\epsilon$-neighborhood multivariate CDF distance introduced in Section 3.3. The empirical CDF is used as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log(n)\right)$.
@@
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log(n)\right)$.
@@
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log(n)\right)$ for S1-S16.
@@ 4.3 Performance metrics
The measurement of the performance of an HDR estimator can be closely related to the specific problem of interest. In this work, for example, our main interest is in detecting abnormal values that, under an HDR framework, would correspond to those points that fall outside the region. This motivates certain metrics of common use in one-class classification problems, which we also employ in this work. However, we emphasize that the derivation of an HDR can in principle have a wider scope compared to a classification goal. Indeed, it would define the region of highest-density points (e.g., normal values), regardless of whether these points have been observed or not. Thus, it would not only allow us to classify an observed point as normal or abnormal, but it would also provide the entire region of normal values, useful, for example, in a prediction setting prior to observing a point.
With that being said, let FP, TP, FN, and TN be the number of false positive, true positive, false negative, and true negative points, respectively, where positive refer to those points that should be outside the true $(1-\alpha)\%$ HDR and negative the others:
@@ 4.4 Simulation results
Comparative performance results are reported in Table 1. Due to space limitation, we focus on four scenarios that are representative of the different copula-induced dependencies and multimodalities, specifically those on the diagonal of Figure 4 (S1, S6, S11, S16), deferring to the Supplementary Material C all the scenarios and sample sizes considered. The Dirichlet case is covered in Section 4.4.1.
@@
In general, copula-based approaches result in the most performing measures, with classification errors (ERR, FPR, FNR) uniformly smaller–or at most equal–than those of the other measures and at no cost in terms of variability. Their advantage is particularly interesting in more complex scenarios, such as the quadrimodal case, where the parametric copula approach (both $M^{\text{PCop}}_{0}$:DE and $M^{\text{PCop}}_{3}$:$\epsilon$-CDF) practically halves the total error rate (ERR) of a standard $M_{0}$:KDE. Although this distance may be considered relatively negligible when looking at the ERR (maximum distance of $0.015$ in S16) and the FPR (maximum distance of $0.008$ in S16), it plays a significant role for the FNR, with an error difference of $0.157$. In practice, the probability of a true positive being missed by the “test” or measure decreases from around $28\%$ ($M_{0}$:KDE; S16) to around $13\%$ ($M^{\text{PCop}}_{0}$:DE, $M^{\text{PCop}}_{3}$:$\epsilon$-CDF; S16). This aspect is well-captured by the alternative performance metrics of F1 and MCC, which offer a more reliable global measure of efficiency compared to the ERR or the Accuracy. As depicted in Table 1–S16, when comparing $M_{0}$:KDE to $M^{\text{PCop}}_{0}$:DE or $M^{\text{PCop}}_{3}$:$\epsilon$-CDF, the MCC, for example, increases from $0.693$ to more than $0.85$. Note that MCC varies from $-1$ (worst value) to $1$ (best value).
@@
When comparing the measures without the copula, no substantial differences can be noticed between $M_{0}$:KDE, $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, with a slightly improved performance of the second. The worst behavior is shown by $M_{2}$:$k$NN-CDF (with FNR being as high as $72\%$ in S16). Note that this was expected as $M_{2}$:$k$NN-CDF follows a univariate rationale, motivating therefore the multivariate $M_{3}$:$\epsilon$-CDF proposal (see Section 3).
@@
Table 1: Performance results in terms of mean (standard deviation) of the compared methods in selected scenarios for sample size $n=500$.

### v2

@@ 3.1 $\boldsymbol{k}$NN-Euclidean distance
This corresponds to the sum of all Euclidean distances of a given point $x$ from its $k$-nearest neighbors. The concept of closeness (“nearest”) is defined according to the Euclidean metric or $L^{2}$-norm denoted by $\lVert\cdot\rVert_{2}$; for $k=1$, the nearest neighbor is the point $x$ itself and, in this case, the distance is zero.
Definition 3 ($k$NN-Euclidean distance).
@@
where $x_{(i)}$ denotes the $i$-th observation in the reordered data such that $0=\lVert x-x_{(1)}\rVert_{2}\leq\dots\leq\lVert x-x_{(i)}\rVert_{2}\leq\dots\lVert x-x_{(n)}\rVert_{2}$.
$M_{1}(x,s_{n},k)$ is entirely based on a metric distance, that is, the Euclidean metric, and represents a sparsity measure. The property follows from the convergence in probability of the $k$-nearest neighbor density estimator (Silverman,, 1996). Similar distances have been used in the literature for nonparametric classification and regression starting from the seminal work of Fix and Hodges, (1951) and Cover and Hart, (1967), which led to the well-known $k$-nearest neighbors ($k$NN) algorithm. Notably, the distance to the nearest neighbors can also be seen as a local density estimate (Loftsgaarden and Quesenberry,, 1965; Silverman,, 1996) and as a special case of a variable-bandwidth KDE with a uniform kernel (Terrell and Scott,, 1992). A large $k$NN distance indicates that the density is small and vice versa. Furthermore, by ranking each data point according to the distance from its $k$-nearest neighbors, this measure can also be used as an outlier score in anomaly detection (Ramaswamy et al.,, 2000).
Compared to other methods, the $k$NN approach has several advantages such as: (i) being purely nonparametric, hence able to flexibly adapt to any continuous distribution; (ii) having a reasonable time complexity; (iii) depending on a unique hyperparameter $k$, whose tuning is relatively simple. The choice of $k$ should be made based on the sample data. Generally, higher values of $k$ reduce the effect of noise; however, they underfit the model–making boundaries between classes less distinct–and are computationally more expensive, especially for large $d$. A general rule of thumb in classification is $k=\lfloor\sqrt{n}\rfloor$, with $n$ the number of samples in the dataset. Further investigation of the role of $k$ in different settings and how it relates to the missclassification error is given in the Supplementary Material B and in Meeker et al., (2017).
@@ 3.2 $\boldsymbol{k}$NN-CDF distance
Although the Euclidean distance and other general notions of measures (including concepts like area or volume) can certainly be useful for capturing relevant insights on the topology of a given set, they may be equipped with a probability measure to better resemble the notion of density. Let $\mathbb{P}$ be a probability measure defined on $\mathbb{R}^{d}$ and denote by $F$ its associated CDF and by $F_{i}$ the CDF of the $i$-th marginal.
@@
The distance $d_{\mathbb{P}}$ can be interpreted as the composition of the (nonlinear) transformation $F\!:\mathbb{R}^{d}\to[0,1]$ and the computation of the ordinary Euclidean distance. It fulfills all the properties of being a proper metric. Furthermore, it has the nice property that the distance between two points is proportional to the probability contained between the two points. However, as defined in Eq. (3), it presents some key limitations for inferring the underlying density and estimating HDRs. Consider for simplicity the univariate Gaussian mixture model illustrated in Figure 2 and the set of points $x_{1}=-2$, $x_{2}=-1.6$, $x_{3}=-0.72$, $x_{4}=0.5$. When evaluating the CDF distance between $x_{2}$ and all other points, we have that $d_{\mathbb{P}}(x_{2},x_{1})\approx d_{\mathbb{P}}(x_{2},x_{3})$, but there is little to say about the density around $x_{1}$ and $x_{3}$, apart from realizing that they are clearly different. One could gain more information if one were to compute the CDF distance from the global mode $x_{1}$, assuming we have this information, noticing that the higher $d_{\mathbb{P}}(x_{1},x_{i})$, for all $i$, the lower the density of $x_{i}$. However, this is no longer verified for multimodal densities. In fact, we have $d_{\mathbb{P}}(x_{1},x_{4})>d_{\mathbb{P}}(x_{1},x_{3})$, but $f(x_{4})>f(x_{3})$, contrasting with the definition of a neighborhood measure.
@@
Motivated by these limitations and inspired by the idea of the $k$NN distance in Definition 3, we propose a new variant consisting of the sum of CDF-distances between a point $x$ and its $k$-nearest neighbors. Here, “nearest” should again be understood according to the Euclidean distance. The rationale is simple: the higher the CDF distances between a point and its neighbors, the higher one expects the density to be at that point. Since the $k$ neighbors of each data point will be at different (Euclidean) distances compared to the $k$ neighbors of the other data points, one needs to properly scale or weight the CDF distances according to this information. The proposed measure is reported in Definition 4.
Definition 4 ($\boldsymbol{k}$NN-CDF distance).
@@
Property 1 (Semimetric).
Consider the $M_{2}$ measure as defined in Definition 4. It can be easily verified that, for any $k\in[1,n]$, if one restricts the CDF distance to $x$ and its $k$-th neighbor $x_{(k)}$, then $M_{2}(x,x_{(k)},k)\!:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a semimetric. In fact, for any set $x,x_{(k)}\in\mathbb{R}^{d}$, it follows that:
@@
The triangular inequality, stating that $M_{2}(x,x_{(k)},k)\leq M_{2}(x,x_{(k^{\prime})},k)+M_{2}(x_{(k^{\prime})},x_{(k)},k)$, for all $x,x_{(k)},x_{(k^{\prime})}\in\mathbb{R}^{d}$, does not hold. It is enough to consider a simple counterexample such as: $d=1$, $F$ the CDF of a standard normal distribution, and three points $x=1<x_{(k)}=2<x_{(k^{\prime})}=3$.
@@ 3.3 $\boldsymbol{\epsilon}$-neighborhood multivariate CDF distance
The measure $M_{2}(x,s_{n},k)$ introduced in Section 3.2 has several advantages; for example: (i) it takes into account the probabilistic information in the data, (ii) it is based on marginal CDFs which may be easy to estimate, and (iii) it has convenient computational times. However, using only the marginal CDFs may compromise their validity whenever the marginal components of a multivariate random variable $X\in\mathbb{R}^{d}$ show a significant dependence structure. In that case, the use of a multivariate CDF may be more appropriate. If the multivariate CDF is unknown, one could use its empirical counterpart. Notice, however, that this estimation part may compromise computational efficiency, especially in high dimensions and for large values of $k$. We therefore focus on an $\epsilon$-neighborhood version, which reduces the number of operations from $k$ to $1$, once $\epsilon$ is determined.
Definition 5 ($\epsilon$-neighborhood multivariate CDF distance).
@@
$\quad-\mathbb{P}(X_{1}\leq x_{1}+\epsilon_{1},X_{2}\leq x_{2}-\epsilon_{2})+\mathbb{P}(X_{1}<x_{1}-\epsilon_{1},X_{2}<x_{2}-\epsilon_{2})$
@@
Property 2 (Density equivalence).
@@ 3.4 Copula-based measures
In summary, one can decompose every $d$-dimensional CDF $F$ or PDF $f$ into a composition of their marginal distribution functions and a $d$-copula. This allows us to redefine both $M_{0}$ in Eq. (2) and $M_{3}$ in Definition 5 in terms of their copula representation. For example, in the case of the $M_{3}$ measure, we may construct its copula-based alternative given in Definition 6.
Definition 6 ($\epsilon$-neighborhood copula-based CDF distance).
@@
$M^{\text{Cop}}_{3}(x,\epsilon)$ $=\frac{\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}C(F_{1}(\nu_{1}),\dots,F_{d}(\nu_{d}))}{\prod_{j=1}^{d}2\epsilon_{j}}\propto\sum_{\nu\in\mathcal{V}}(-1)^{n(\nu)}C(F_{1}(\nu_{1}),\dots,F_{d}(\nu_{d})),$ (7)
@@
The main advantage of this representation over the one involving the joint CDF is that the estimation of the multivariate distribution–when unknown–is performed through the estimation of the (univariate) marginals, evading thus the curse of dimensionality (see e.g., Nagler and Czado,, 2016). Furthermore, copulae offer a flexible framework that captures complex dependence structures while offering direct control over the marginals.
@@ 4.1 Simulation setup
We now evaluate and compare the proposed measures in an extensive number of simulated bivariate settings that vary according to the complexity of the data. In particular, we consider scenarios with different dependence structures induced by the copula model (e.g., tail or asymmetric dependencies), different marginal distributions (e.g., heavy-tailed or multimodal distributions), as well as different sample sizes.
@@
    We finally cover different sample sizes, with $n\in\{50,100,500,1000\}$.
@@
By exploring such an extended number of simulation setups, we hope to cover a broad spectrum of potential scenarios that may occur in practice. In addition to offering a comprehensive understanding of the performances of the different measures, it may provide guidance to applied scientists who may need to choose the most ideal measure for their specific application. For example, the problem of doping detection when relying on the two primary biomarkers of the hematological module could well be related to scenario S15 (Gaussian and Gaussian mixture marginals and a Clayton copula; see Deliu and Liseo, (2024) for an illustrative example).
@@ 4.2 Evaluated measures
This represents the $\epsilon$-neighborhood multivariate CDF distance introduced in Section 3.3. The empirical CDF is used as the CDF estimator. The optimal choice of the hyperparameter $\epsilon$ is based on an extensive cross-validation procedure, leading to the following heuristic for S1-S16 as a result of an exponential decay model fit with respect to the sample size: $\epsilon=\exp\left(2.13-0.3\log{n}\right)$.
@@
Fully nonparametric indirect estimation of the $\epsilon$-CDF measure with copula. We use the empirical CDF for estimating the univariate marginals, and KDE with the transformation local likelihood estimator and nearest-neighbor bandwidth for the copula density. The optimal choice of the hyperparameter $\epsilon$ follows the same strategy as the previous measure, leading to the following heuristic for S1-S16: $\epsilon=\exp\left(1.74-0.26\log{n}\right)$.
@@
Parametric indirect CDF estimation with copula. We adopt a fully parametric approach (with maximum likelihood fitting) to estimate both marginals and the copula. For the copula model, we select the best model using the AIC criterion; no misspecification is introduced for the marginals. The optimal choice of the hyperparameter $\epsilon$ follows the heuristic $\epsilon=\exp\left(1.60-0.41\log{n}\right)$ for S1-S16.
@@ 4.3 Performance metrics
The measurement of the performance of an HDR estimator can be closely related to the specific problem of interest. In a context where the interest is in detecting abnormal values, for example, estimating an HDR would allow to understand which points fall outside the normal-points region. This motivates certain metrics of common use in one-class classification problems, which we also employ in this work. However, we emphasize that the derivation of an HDR can in principle have a wider scope compared to a classification goal. Indeed, it would define the region of highest-density points (e.g., normal values), regardless of whether these points have been observed or not. Thus, it would not only allow us to classify an observed point as normal or abnormal, but it would also provide the entire region of normal values, useful, for example, in a prediction setting prior to observing a point.
Let FP, TP, FN, and TN be the number of false positive, true positive, false negative, and true negative points, respectively, where positive refers to those points that should be outside the true $(1-\alpha)\%$ HDR and negative the others:
@@ 4.4 Simulation results
Comparative performance results are reported in Table 1 in terms of their mean (standard deviation) across a number of $1000$ independent Monte Carlo (MC) replicates. We primarily focus on discussing four scenarios that are representative of the different copula-induced dependencies and multimodalities, specifically those on the diagonal of Figure 3 (S1, S6, S11, S16), and the sample size $n=500$; all the other scenarios and sample sizes are deferred to the Supplementary Material C. The Dirichlet case is covered in Section 4.4.1.
@@
In general, copula-based approaches result in the most performing measures, with classification errors (ERR, FPR, FNR) uniformly smaller than those of the other measures and at no cost in terms of variability. Their advantage is particularly interesting in more complex scenarios, such as the quadrimodal case, where the parametric copula approach (both $M^{\text{PCop}}_{0}$:DE and $M^{\text{PCop}}_{3}$:$\epsilon$-CDF) practically halves the total error rate (ERR) of a standard $M_{0}$:KDE. Although the difference may be considered relatively negligible when looking at the ERR (maximum difference of $0.015$ in S16) and the FPR (maximum distance of $0.008$ in S16), it plays a significant role for the FNR, with an error difference of $0.157$. In practice, the probability of a true positive being missed by the “test” or measure decreases from around $28\%$ ($M_{0}$:KDE; S16) to around $13\%$ ($M^{\text{PCop}}_{0}$:DE, $M^{\text{PCop}}_{3}$:$\epsilon$-CDF; S16). This aspect is well-captured by the alternative performance metrics of F1 and MCC, which offer a more reliable global measure of efficiency compared to the ERR or the Accuracy. As shown in Table 1–S16, when comparing $M_{0}$:KDE to $M^{\text{PCop}}_{0}$:DE or $M^{\text{PCop}}_{3}$:$\epsilon$-CDF, the MCC, for example, increases from $0.693$ to more than $0.85$. Note that MCC varies from $-1$ (worst value) to $1$ (best value).
When comparing the non-copula based measures, no substantial difference is noticed between $M_{0}$:KDE, $M_{1}$:$k$NN-Eucl and $M_{3}$:$\epsilon$-CDF, with a slightly improved performance of the second. In particular, $M_{1}$:$k$NN-Eucl provides the only exception to the uniform advantage of copula-based approaches. This occurs in S1 for $n=50$, where $M_{1}$:$k$NN-Eucl shows an enhancement, although negligible (results are given in the Supplementary material C; Table 2 and Figure 31). Interestingly, S1 represents a scenario with heavy-tailed marginals, i.e., Student-$t$ distribution with $\nu=2$, which translates into a substantially wider $95\%$ HDR (see Figure 3). Relatively good results are also achieved in the other heavy-tailed cases (S5, S9, and S13), and only in small samples (in particular $n=50$; see Figure 31 in the Supplementary material). Outside these scenarios, compared to the classical $M_{0}$:KDE, $M_{1}$:$k$NN-Eucl has typically superior performances in more complex cases, in particular, in all quadrimodal distributions (S4, S8, S12, S16), and for smaller sample sizes ($n=50$ and $n=100$). As the sample size increases ($n>100$), $M_{0}$:KDE shows some improvements, especially in simpler settings with Gaussian marginals (result are given in the Supplementary material C).
@@
The worst behavior is shown by $M_{2}$:$k$NN-CDF (with FNR being as high as $72\%$ in S16). Note that this was expected as $M_{2}$:$k$NN-CDF follows a univariate rationale, motivating therefore the multivariate $M_{3}$:$\epsilon$-CDF proposal (see Section 3).
@@
Table 1: Performance results of the compared methods in selected scenarios for sample size $n=500$. Data are summarized in terms of mean (standard deviation) across $1000$ independent Monte Carlo (MC) replicates.